OpenStudy (anonymous):

( sqrt(6-x) -2 ) * ( sqrt(3-x) + 1) ??

OpenStudy (anonymous):

(√6-x-2)∙(√3-x+1) is this what the equation looks like?

OpenStudy (anonymous):

only the 6-x is squarerooted. same with the 3-x

OpenStudy (anonymous):

us the equation editor in windows to show me the problem

OpenStudy (anonymous):

\[(\sqrt{6-x}-2) \times ( \sqrt{3-x} + 1) \]

OpenStudy (anonymous):

ok that helps.

OpenStudy (anonymous):

now, do you want to know how to work it out the right way or just the answer with no clue how you got it.?

OpenStudy (anonymous):

i want to know how to work it out :) thank you

OpenStudy (anonymous):

(~(6-x)-2)*(~(3-x)+1) Reorder the polynomial 6-x alphabetically from left to right, starting with the highest order term. (~(-x+6)-2)*(~(3-x)+1) Reorder the polynomial 3-x alphabetically from left to right, starting with the highest order term. (~(-x+6)-2)*(~(-x+3)+1) Multiply each term in the first group by each term in the second group using the FOIL method. FOIL stands for First Outer Inner Last, and is a method of multiplying two binomials. First, multiply the first two terms in each binomial group. Next, multiply the outer terms in each group, followed by the inner terms. Finally, multiply the last two terms in each group. (~(-x+6)*~(-x+3)+~(-x+6)*1-2*~(-x+3)-2*1) Simplify the FOIL expression by multiplying and combining all like terms. (~(-x+6)~(-x+3)+~(-x+6)-2~(-x+3)-2) Remove the parentheses around the expression ~(-x+6)~(-x+3)+~(-x+6)-2~(-x+3)-2. ~(-x+6)~(-x+3)+~(-x+6)-2~(-x+3)-2 Multiply ~(-x+6) by ~(-x+3) to get ~(-x+6*-x+3). ~(-x+6*-x+3)+~(-x+6)-2~(-x+3)-2 Multiply 6 by -x to get -6x. ~(-x-6x+3)+~(-x+6)-2~(-x+3)-2 Since -x and -6x are like terms, subtract 6x from -x to get -7x. ~(-7x+3)+~(-x+6)-2~(-x+3)-2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

how would I factor that final answer by 2-x?

OpenStudy (anonymous):

nevermind i got it :)

OpenStudy (anonymous):

let me know what you got, no one has ever asked that before

OpenStudy (anonymous):

http://openstudy.com/groups/mathematics/updates/4e8e488c0b8b543d4218a62b i followed those steps :)

OpenStudy (anonymous):

it was just a different method. thank you very much though!

OpenStudy (anonymous):

your very welcome, I guess I just like the loooooooong route, lol